OpenStudy (anonymous):

Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents (10x^-5y^0/20x^2y^-4)^-4

OpenStudy (anonymous):

what was your answer? i'll tell you whether you did it correctly or not. you need to remember that (a^b)^c = a^(bc) and a^-b = 1/(a^b)

OpenStudy (anonymous):

AdequateDisclosure , your answer is incorrect. try again.

OpenStudy (anonymous):

let me view it again

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):